The projected worth (in millions of dollars) of a large company is modeled by the equation y = 206(1.1) x. The variable x represents the number of years since 2000. What is the projected annual percent of growth, and what should the company be worth in 2011?

Answer by stanbon(75887)

The projected worth (in millions of dollars) of a large company is modeled by the equation y = 206(1.1)^x. The variable x represents the number of years since 2000. What is the projected annual percent of growth, and what should the company be worth in 2011?
-----
y = 206*(1+0.10)^x
Percent growth:: 0.10 = 10%
----
f(2) = 206(1.1)^2 = 249.26 million
--------------
Cheers,
Stan H.
------------